OpenStudy (anonymous):

A tank contains 1000 L of brine (that is, salt water) with 15 kg of dissolved salt. Pure water enters the top of the tank at a constant rate of 10 L / min. The solution is thoroughly mixed and drains from the bottom of the tank at the same rate so that the volume of liquid in the tank is constant. Find a differential equation expressing the rate at which salt leaves the tank.

OpenStudy (anonymous):

The answer is\[\frac{ds}{dt}=-\frac{s}{100}\]

OpenStudy (anonymous):

I don't get why the rate of change depends on the amount of salt in the tank.

OpenStudy (ranga):

Let "s" kg of salt be in the tank at time "t". At delta_t minutes later, delta_s kg of salt would have left the tank and so the amount of salt in the tank will be less by delta_s. The concentration of salt in the tank at time t is: s / 1000 kg/liter In delta_t minutes, 10 * delta_t liters of liquid would have left the tank and the amount of salt that leaves with it = 10 * delta_t * s / 1000 kg. = delta_t * s / 100 delta_s = - delta_t * s / 100 ds/dt = -s / 100

OpenStudy (anonymous):

Thanks, for some reason I didn't realize that the salt would spread into the pure water too which means that there would be less and less salt leaving over time.